# Omega: Efficient Occlusion-Aware Navigation for Air-Ground Robots in Dynamic Environments via State Space Model

### v1

@@ Abstract
Air-ground robots (AGRs) are widely used in surveillance and disaster response due to their exceptional mobility and versatility (i.e., flying and driving). Current AGR navigation systems perform well in static occlusion-prone environments (e.g., indoors) by using 3D semantic occupancy networks to predict occlusions for complete local mapping and then computing Euclidean Signed Distance Field (ESDF) for path planning. However, these systems face challenges in dynamic scenes (e.g., crowds) due to limitations in perception networks’ low prediction accuracy and path planners’ high computation overhead.
In this paper, we propose Omega, which contains OccMamba with an Efficient AGR-plAnner to address the above-mentioned problems. OccMamba adopts a novel architecture that separates semantic and occupancy prediction into independent branches, incorporating the mamba module to efficiently extract semantic and geometric features in 3D environments. This ensures the network can learn long-distance dependencies and improve prediction accuracy. Features are then combined within the Bird’s Eye View (BEV) space to minimise computational overhead during feature fusion. The resulting semantic occupancy map is integrated into the local map, providing occlusion awareness of the dynamic environment. Our AGR-Planner utilizes this local map and employs Kinodynamic A* search and gradient-based trajectory optimization for ESDF-free and energy-efficient planning. Experiments demonstrate that OccMamba outperforms the state-of-the-art 3D semantic occupancy network with 25.0% mIoU. End-to-end navigation experiments in dynamic scenes verify OMEGA’s efficiency, achieving a 96% average planning success rate. Code and Video Page: https://jmwang0117.github.io/OMEGA/.
@@ I INTRODUCTION
Unfortunately, while these methods have shown success in static occlusion-prone environments (e.g., indoor or forest), they struggle in dynamic environments with moving obstacles (Fig. 1). This limitation can be attributed to two main factors: the perception network’s low prediction accuracy and high inference latency, and the path planner’s high computational overhead. Firstly, although the use of 3D convolutions [5] and transformers [9, 10] in 3D semantic occupancy networks has improved occupancy prediction accuracy, it has also introduced computational overhead. This increased computational demand challenges real-time inference on resource-constrained AGR devices (e.g., Jetson Xavier NX). Secondly, current AGR path planners [2, 4] allocate nearly 70% of the total local planning time to construct the Euclidean Signed Distance Field (ESDF) map. This extensive processing time is impractical for dynamic environments with rapidly changing scenes, as it heightens the risk of collisions (Table I).
Our key insight to address the above limitations is to propose an efficient 3D semantic occupancy network that ensures real-time inference while improving prediction accuracy. We draw inspiration from state space models (SSMs) [11] and their improved versions, such as Mamba [12] and Mamba-2 [13] networks. By integrating mamba blocks into a new 3D semantic occupancy network architecture, we aim to enable the network to model long-distance dependencies and perform parallel feature learning. This approach facilitates real-time occlusion prediction in highly dynamic environments, resulting in more complete local maps. Furthermore, while Zhou et al. [14] developed an ESDF-free path planner for quadcopters, it does not address AGR-specific requirements in dynamic scenarios, particularly energy efficiency and dynamic constraints. Consequently, exploring ESDF-free AGR path planners is crucial to improving overall navigation performance.
@@ II-A Dynamic Navigation System for AGRs
Numerous researchers have explored various aerial-ground robot configurations, such as incorporating passive wheels [2, 15, 16], cylindrical cages [17], or multi-limb [18] onto drones. Recently, Fan et al. [1] address ground-aerial motion planning. Their approach initially employs the A* algorithm to search for a geometric path as guidance, favouring ground paths by adding extra energy costs to aerial paths. Zhang et al. [2] proposed a path planner and controller capable of path searching, but it relies on an ESDF map. The intensive computation and limited perception of occluded areas lead to a low success rate in path planning and increased energy consumption. Wang et al. [4] proposed AGRNav, the first AGR navigation system with occlusion perception capability. Although it performs well in static environments, its simple perception network structure design and the defects of the path planner make it difficult to operate in complex and changing dynamic environments. Our work aims to explore new AGR navigation systems for efficient navigation in dynamic environments.
@@ IV-B Gradient-Based B-spline Trajectory Optimization
Collision Avoidance Force Estimation: Inspired by [14], for each control point on the collision trajectory segment, vector $v$ (i.e., a safe direction pointing from inside to outside of that obstacle) is generated from $\iota$ to $\tau$ and $p$ is defined at the obstacle surface (in Fig. 4a). With generated $\left\{p,v\right\}$ pairs, the planner maximizes $D_{ij}$ and returns an optimized trajectory. The obstacle distance $D_{ij}$ if $i^{th}$ control point $Q_{i}$ to $j^{th}$ obstacle is defined as:
@@
Because the guide path $\tau$ is energy-saving, the generated path is also energy efficient (in Fig. 4a). To discover multiple viable paths that navigate through different local minima in a dynamic environment, we apply the method described in [14]. This method generates distance fields in various directions by reversing vector $v_{1}$ to obtain $v_{2}=-v_{1}$. Following this, a search process identifies a new anchor point $p_{2}$ on the surface of the obstacle along $v_{2}$, as shown in Fig. 4a. This approach enables us to evaluate alternative trajectories and select the most cost-effective path for navigation.
@@ V-A Evaluation setup
3D Semantic Occupancy Network: We trained OccMamba using the outdoor SemanticKITTI dataset on a single NVIDIA 3090 GPU. This dataset [33] focuses on semantic occupancy prediction with LiDAR point clouds and camera images. Specifically, the ground truth semantic occupancy is represented as the $[256,256,32]$ voxel grids. Each voxel is $[0.2m,0.2m,0.2m]$ large. OccMamba was trained for 80 epochs with a batch size of 6, using the AdamW optimizer [36] at an initial learning rate of 0.001, and input point cloud augmentation by random flipping along the $x-y$ axis.
@@
Real-world Experiment: In our real-world experiments, Omega was operational on a customed AGR platform, illustrated in Fig. 2. The platform employed a RealSense D435i for point cloud acquisition, a T265 camera for visual-inertial odometry, and a Jetson Xavier NX for real-time onboard computation. OccMamba was optimized with TensorRT for efficient offline inference, guaranteeing smooth and responsive navigation. Our evaluation is around the following key questions:
RQ1: How does OccMamba’s efficiency compare to baseline models? (§ V-B)
RQ2: How well does Omega navigate in dynamic simulation environments? (§ V-C)
RQ3: How effectively does Omega navigate in five real-world dynamic scenarios? (§ V-D)
@@ V-D Real-world Air-Ground Robot Navigation
We demonstrated OMEGA’s superior efficiency and energy conservation in 5 dynamic scenarios with a velocity cap of 1.5 m/s. Each navigation method was tested 10 times per scenario for reliable performance comparison. We achieved a notable 96% average success rate (Fig. 7a) in 5 dynamic scenarios with OMEGA, exhibiting lower average energy consumption relative to competitors (Fig. 7c). Specifically, in scenarios B and C, OMEGA recorded energy reductions of 22.1% and 17.28%, respectively, against AGRNav. These improvements are largely due to our OccMamba module’s swift computational ability to predict obstacle distributions in non-visible areas, enabling the AGR system to avoid potential impediments. When integrated with the AGR-Planner, this foresight allows for the rapid generation of multiple candidate paths, optimizing for both energy efficiency and reduced traversal time (Fig. 7b).
@@ VI CONCLUSION
In this letter, we introduce OMEGA, an advanced system for air-ground robot (AGR) navigation in dynamic environments. The system includes OccMamba, which efficiently separates semantic prediction and scene completion, using mamba blocks for fast and precise real-time semantic occupancy mapping. This creates a comprehensive local map without occlusions. Based on the local map, the AGR-Planner component of OMEGA generates safe, smooth, and dynamically feasible trajectories. Our extensive testing shows that OccMamba outperforms state-of-the-art methods with a 59.9% IoU at 22.1 FPS. When integrated into the OMEGA system, the combination of OccMamba and AGR-Planner results in an impressive 96% average success rate for planning in real-world dynamic scenarios.
VII Acknowledgments
We thank all anonymous reviewers for their helpful comments.

### v2

@@ Abstract
Air-ground robots (AGRs) are widely used in surveillance and disaster response due to their exceptional mobility and versatility (i.e., flying and driving). Current AGR navigation systems perform well in static occlusion-prone environments (e.g., indoors) by using 3D semantic occupancy networks to predict occlusions for complete local mapping and then computing Euclidean Signed Distance Field (ESDF) for path planning. However, these systems face challenges in dynamic scenes (e.g., crowds) due to limitations in perception networks’ low prediction accuracy and path planners’ high computation overhead. In this paper, we propose Omega, which contains OccMamba with an Efficient AGR-plAnner to address the above-mentioned problems. OccMamba adopts a novel architecture that separates semantic and occupancy prediction into independent branches, incorporating the mamba module to efficiently extract semantic and geometric features in 3D environments. This ensures the network can learn long-distance dependencies and improve prediction accuracy. Features are then combined within the Bird’s Eye View (BEV) space to minimise computational overhead during feature fusion. The resulting semantic occupancy map is integrated into the local map, providing occlusion awareness of the dynamic environment. Our AGR-Planner utilizes this local map and employs Kinodynamic A* search and gradient-based trajectory optimization for ESDF-free and energy-efficient planning. Experiments demonstrate that OccMamba outperforms the state-of-the-art 3D semantic occupancy network with 25.0% mIoU. End-to-end navigation experiments in dynamic scenes verify OMEGA’s efficiency, achieving a 96% average planning success rate.
@@ I INTRODUCTION
Unfortunately, while these methods have shown success in static occlusion-prone environments (e.g., indoor or forest), they struggle in dynamic environments with moving obstacles (Fig. 1). This limitation can be attributed to two main factors: the perception network’s high inference latency and low prediction accuracy, and the path planner’s high computational overhead. Firstly, despite the improved accuracy of transformer-based 3D semantic occupancy prediction networks [9, 10] in simultaneously predicting static and dynamic obstacles for occlusion-free local mapping, their high computational overhead hinders deployment on resource-constrained AGR devices (e.g., Jetson Xavier NX). Secondly, current AGR path planners [2, 4] allocate nearly 70% of the total local planning time to construct the Euclidean Signed Distance Field (ESDF) map. This extensive processing time is impractical for dynamic environments with rapidly changing scenes, as it heightens the risk of collisions (Table I).
Our key insight to address the above limitations is to propose an efficient 3D semantic occupancy network that ensures real-time inference while improving prediction accuracy. We draw inspiration from state space models (SSMs) [11] and their improved versions, such as Mamba [12, 13]. By integrating mamba blocks into a new 3D semantic occupancy network architecture, we aim to enable the network to model long-distance dependencies and perform parallel feature learning. This approach facilitates real-time occlusion prediction in highly dynamic environments, resulting in more complete local maps. Furthermore, while Zhou et al. [14] developed an ESDF-free path planner for quadcopters, it does not address AGR-specific requirements in dynamic scenarios, particularly energy efficiency and dynamic constraints. Consequently, exploring ESDF-free AGR path planners is crucial to improving overall navigation performance.
@@ II-A Dynamic Navigation System for AGRs
Researchers have explored various aerial-ground robot configurations, such as incorporating passive wheels [2, 16, 17], cylindrical cages [18], or multi-limb [19] onto drones. Recently, Fan et al. [1] address ground-aerial motion planning. Their approach initially employs the A* algorithm to search for a geometric path as guidance, favouring ground paths by adding extra energy costs to aerial paths. Zhang et al. [2] proposed a path planner and controller capable of path searching, but it relies on an ESDF map. The intensive computation and limited perception of occluded areas lead to a low success rate in path planning and increased energy consumption. Wang et al. [4] proposed AGRNav, the first AGR navigation system with occlusion perception capability. Although it performs well in static environments, its simple perception network structure design and the defects of the path planner make it difficult to operate in complex and changing dynamic environments. Our work aims to explore new AGR navigation systems for efficient navigation in dynamic environments.
@@ IV-B Gradient-Based B-spline Trajectory Optimization
Collision Avoidance Force Estimation: For each control point on the collision trajectory segment, vector $v$ (i.e., a safe direction pointing from inside to outside of that obstacle) is generated from $\iota$ to $\tau$ and $p$ is defined at the obstacle surface (in Fig. 4a). With generated $\left\{p,v\right\}$ pairs, the planner maximizes $D_{ij}$ and returns an optimized trajectory. The obstacle distance $D_{ij}$ if $i^{th}$ control point $Q_{i}$ to $j^{th}$ obstacle is defined as:
@@
Because the guide path $\tau$ is energy-saving, the generated path is also energy efficient (in Fig. 4a). Inspired by [14], we discover multiple viable paths that navigate through different local minima in a dynamic environment. This is achieved by generating distance fields in various directions through reversing vector $v_{1}$ to obtain $v_{2}=-v_{1}$. Following this, a search process identifies a new anchor point $p_{2}$ on the obstacle’s surface along $v_{2}$, as shown in Fig. 4a. This approach enables us to evaluate alternative trajectories and select the most cost-effective path for navigation.
@@ V-A Evaluation setup
3D Semantic Occupancy Network: We trained OccMamba on the outdoor SemanticKITTI dataset [34] using a single NVIDIA 3090 GPU. The dataset provides LiDAR point clouds for semantic occupancy prediction, with ground truth represented as $[256,256,32]$ voxel grids (each voxel measuring $0.2m^{3}$). Training spanned 80 epochs with a batch size of 6, utilizing the AdamW optimizer [37] with an initial learning rate of 0.001. We applied random x-y axis flipping for input point cloud augmentation. The total training time was approximately 40 hours.
@@
Real-world Experiment: In our real-world experiments, Omega was deployed on a custom AGR platform (Fig. 2) equipped with a RealSense D435i for point cloud acquisition, a T265 camera for visual-inertial odometry, and a for real-time onboard computation. The best-performing pre-trained OccMamba model (IoU=59.9) was deployed offline, using the same settings from AGRNav [4] . Optimized with TensorRT on Jetson Xavier NX, it efficiently completed local maps for smooth navigation.
@@ V-D Real-world Air-Ground Robot Navigation
We demonstrated OMEGA’s superior efficiency and energy conservation in 5 dynamic scenarios with a velocity cap of 1.5 m/s. Each navigation method was tested 10 times per scenario for reliable performance comparison. We achieved a notable 96% average success rate (Fig. 7a) in 5 dynamic scenarios with OMEGA, exhibiting lower average energy consumption relative to competitors (Fig. 7c). Specifically, in scenarios B and C, OMEGA recorded energy reductions of 22.1% and 17.28%, respectively, against AGRNav. These improvements are largely due to our OccMamba module’s swift computational ability to predict obstacle distributions in non-visible areas, enabling the AGR system to avoid potential impediments. When integrated with the AGR-Planner, this foresight allows for the rapid generation of multiple candidate paths, optimizing for both energy efficiency and reduced traversal time (Fig. 7b). Please zoom in on Fig. 7 to see more detailed qualitative and quantitative results.
@@ VI CONCLUSION
In this letter, we introduce OMEGA, an advanced system for air-ground robot (AGR) navigation in dynamic environments. OMEGA features OccMamba, which efficiently generates comprehensive, occlusion-free local maps using mamba blocks for real-time semantic occupancy mapping. The AGR-Planner component then utilizes these maps to produce safe and dynamically feasible trajectories. Our evaluations show OccMamba outperforming state-of-the-art methods with 59.9% IoU at 22.1 FPS. When integrated into OMEGA, the system achieves a 96% average success rate in real-world dynamic scenarios.